OpenStudy (anonymous):

solve: the square of 6g = (j+y) for g.

OpenStudy (anonymous):

\[solve: \sqrt{6g}=\left( j+y \right) for g\]

OpenStudy (cwrw238):

square root right?

OpenStudy (anonymous):

yes, i just added the equation.

OpenStudy (cwrw238):

square both sides 6g = (j + y)^2 now its easy

OpenStudy (cwrw238):

divide both sides by 6

OpenStudy (anonymous):

when i did that, i came up with g= (j+y)^2 / 6,,, but the answer is g= (j^2 +2jy+y^2) / 6 I don't get how to get the correct answer

OpenStudy (cwrw238):

ah - thats same thing - they've expanded (j + y)^2 which comes to ( j^2 + 2jy + y^2)

OpenStudy (cwrw238):

so answer is ( j^2 + 2jy + y^2) / 6

OpenStudy (anonymous):

oh i see now,, tyyt!!!